Circular No. 8034

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

2001 QC_{298}

K. Noll and D. Stephens, Space Telescope Science Institute; W. Grundy, Lowell Observatory; D. Cruikshank, Ames Research Center, NASA; S. Tegler, Northern Arizona University; and W. Romanishin, University of Oklahoma, report the detection of a binary companion to the transneptunian object (TNO) 2001 QC_{298} (MPEC 2002-P49, 2001-T40), from observations made on Oct. 4.646–4.672 UT with camera 2 of the Near-Infrared Camera Multi-Object Spectrograph on the Hubble Space Telescope. The TNO was observed in the F110W and F160W filters (yielding ≈ J and H bands), with two exposures in each filter. The object is extended in all four images. Contour and surface plots identify two components of the binary, separated by an angular distance of 0′′.17 ± 0′′.08 at position angle 355^{o} ± 10^{o}. The spacecraft was tracking 2001 QC_{298} during the observations. Stars and galaxies moved by 4.3 pixels between pairs of exposures in a direction roughly perpendicular to the extension in the TNO image, while 2001 QC_{298} remained at the same position to within 0.1 pixel. A geocentric distance of 39.72 AU at the time of observation yields an apparent separation of the objects in the plane of the sky of 5000 ± 2000 km.

SUPERNOVAE 2002jq–2002jx P. Challis, on behalf of the ESSENCE project (cf. IAUC 8021), reports

the discovery of eight supernovae, each present in at least two bands (V, R, I) taken with the Blanco 4-m telescope (+ 8K Mosaic II imager) at Cerro Tololo. The supernovae were identified via subtraction of images taken on Oct. 8–14 UT (when none of the new objects was visible, to limiting mag R ∼ 23.5) from those taken on the tabulated discovery dates, with further confirming images taken on Dec. 8, 12, 14, and 16, and with spectroscopy obtained on Nov. 28–Dec. 8 by project members at the Gemini (+ GMOS), VLT (+ FORS1), and Magellan (+ LDSS2) telescopes.

SN 2002 UT α_{2000} δ_{2000} R z Type
2002jq Nov. 5 23^{h}35^{m}57.^{s}96 −10^{o}05′56′′.8 21.9 0.5 Ia
2002jr Nov. 9 2 04 41.03 − 5 09 40.7 22.7 0.4 Ia
2002js Nov. 9 2 20 35.39 − 9 34 43.9 22.9 0.4 Ia
2002jt Nov. 28 0 13 36.70 −10 08 24.0 22.6 0.5 Ia
2002ju Dec. 2 2 20 11.00 − 9 04 37.5 21.6 0.3 Ia
2002jv Dec. 6 2 04 34.82 − 3 51 57.9 22.6 0.2 II
2002jw Dec. 6 2 30 00.52 − 8 36 22.4 22.8 0.4 Ia
2002jx Dec. 6 0 26 54.13 + 0 22 49.7 23.5 0.2 II
2002 December 17  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green